OpenStudy (anonymous):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2 h, and Car B traveled the distance in 1.5 h. Car B traveled 15 mph faster than Car A. How fast did Car B travel? mph

OpenStudy (damonte212):

Let speed of car B be x Speed of car A is (x - 15) Distance traveled by car B in 1.5 h = x * 1.5 Distance traveled by car A in 2 h = (x - 15) * 2 These two distances are equal Therefore, 1.5x = 2(x - 15) 1.5x = 2x - 30 subtract 1,5x 0 = 0.5x - 30 add 30 30 = 0.5x divide by 0.5 30/0.5 = x x = 60 60 is the Anawer

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

could you help with one more

OpenStudy (damonte212):

NP

OpenStudy (damonte212):

yes

OpenStudy (anonymous):

Francine and Cheryl received equal scores on a test made up of multiple choice questions and an essay. Francine got 34 multiple choice questions correct and received 14 points for her essay. Cheryl got 30 multiple choice questions correct and received 22 points for her essay. How many points was each multiple choice question worth? points I can't seem to figure problems out like these

OpenStudy (damonte212):

34m + 14 = 30m + 22 4m = 8 m = 2

OpenStudy (anonymous):

So the answer is 2points

OpenStudy (damonte212):

Right

OpenStudy (anonymous):

Ok thanks,what grade are you?

OpenStudy (damonte212):

Sophmore

OpenStudy (anonymous):

one more?

OpenStudy (damonte212):

Sure lol

OpenStudy (anonymous):

You have no idea how much your helping lol In March, Isaiah and Freddie had the same amount of money in their savings accounts. In April, Isaiah deposited $210 into his account. Freddie increased the money in his account by 15%. When they compared their balances, they found that they were still equal. How much money did they both have in their accounts in March?

OpenStudy (damonte212):

x + 210 = 1.15x 210 = .15x 1400 = x

OpenStudy (anonymous):

1400? the answer

OpenStudy (damonte212):

Right